OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x-4y=-24 x+y=-1

OpenStudy (anonymous):

X=-4, y=3

OpenStudy (anonymous):

3x-4y=-24 x+y=-1/*4 3x-4y=-24 4x+4y=-4 7x=-28 x=-4 y= 3